Your company has sent you on an overseas assignment in Budapest, Hungary. The current indirect rate for the Hungarian forint is 187.90. If the 18-month interest rate in the U.S. is 3 percent and the 18-month interest rate in Hungary is 8.5% (annually), what is the 18-month forward rate?
Answer #1

18 month interest rate in US =3%
18 Month interest in Hungary =8.5%
Number of years =18/12 =1.5
18 month forward rate =Indirect rate*(1+US rate)/(1+Hungary rate) =187.90*((1+3%)/(1+8.5%))^1.5 =173.80